# Learning UAV-based path planning for efficient localization of objects using prior knowledge

### v1

@@ 5 Conclusion
In this study, we showed that a learned search policy can increase the efficiency of finding objects using a UAV, particularly when there was some prior knowledge available. When the objects were non-uniformly distributed, the learned search policy outperformed the baseline row-by-row flight path. The learned policy was robust against errors in the detection, and different qualities in prior knowledge only had minor influence on the number of found objects. In addition, the RL agent was able to learn when the search task had to be terminated and when it was not profitable to continue searching.
@@
In conclusion, learning a search policy improves the efficiency of a search task for UAVs for applications where the target objects, such as weeds and disease infections, are non-uniformly distributed and when there is some level of prior knowledge. Learning to search objects in the absence of prior knowledge remains a topic for future work.
@@
The simulation and RL network is made available on https://github.com/wur-abe/rl_drone_object_search.

### v2

@@ 5 Conclusion
In this study, we showed that a learned search policy can increase the efficiency of finding weeds using a UAV, particularly when there was some prior knowledge available. When the weeds were non-uniformly distributed, the learned search policy outperformed the baseline row-by-row flight path. The learned policy was robust against errors in the detection, and different qualities in prior knowledge only had a minor influence on the number of weeds found. In addition, the RL agent was able to learn when the search task had to be terminated and when it was not profitable to continue searching. Finally, we demonstrated the transferability of the learned policy to real-world data, achieving a 66% shorter flight path compared to a row-by-row flight path at the cost of a 10% lower percentage of found weeds.
@@
In conclusion, learning a search policy improves the efficiency of a search task for UAVs for applications where the target objects, such as weeds, are non-uniformly distributed and when there is some level of prior knowledge. Learning to search these objects in the absence of prior knowledge and evaluation of the RL policy on a real drone remains a topic for future work.
@@
The simulation and the RL network are made available on https://github.com/wur-abe/rl_drone_object_search.